Circular No. 8298

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

POSSIBLE SUPERNOVA ASSOCIATED WITH Q0957+561 C. Y. Peng and C. D. Impey, University of Arizona; H.-W. Rix, Max-

Planck-Institut für Astronomie, Heidelberg; C. S. Kochanek, Ohio State University; E. E. Falco, J. Lehar, and B. A. McLeod, Harvard-Smithsonian Center for Astrophysics; and C. R. Keeton, University of Chicago, report the discovery, on 28 images taken over 9 hr on Feb. 4 with the Hubble Space Telescope NICMOS camera 2, of a possible supernova (infrared mag H = 22.9 ± 0.1) in the gravitational-lens system Q0957+561. The new object is located ∼ 1′′.10 east and 0′′.08 north from image ‘B’ of the quasar (α = 10^{h}01^{m}20^{s}.8, δ = +55^{o}53′50′′, equinox 2000.0). A similar frame taken on 1998 May 30 shows no stellar object at this location (limiting mag H = 24). The new object may be located in the outskirts of the lensing galaxy (z = 0.36) or in the lensed host galaxy of the quasar (z = 1.41). If the possible supernova is lensed, its intrinsic magnitude is roughly 22.4 after correcting the lens magnification, or M_{R} ∼ −20.4. The new object appears near the trailing image B of the quasar; thus, while one would expect a counter image of the possible supernova to appear near the leading ‘A’ image, it would be 173 days (rest frame) further along the light curve and would be too faint to detect. And, in fact, one sees no counterpart near the ‘A’ image.

TRANSITS OF MARS I AND II J. Bell, Cornell University; M. Lemmon, Texas A&M University; and

M. Wolff, Space Science Institute; and the Mars Exploration Rover ‘Athena’ Science Team, report the imaging of transits of Mars I (Phobos) and Mars II (Deimos) across the sun with the Pancam 1-cm f/20 binocular CCD camera (+ solar neutral-density 5 filter; passband 880 ± 20 nm; exposure lengths 0.06 and 0.10 s, taken at 10-s intervals) onboard the ‘Opportunity’ Rover craft located on the planet’s surface at longitude 354^{o}.47417 east and latitude 1^{o}.94828 south. This is apparently the first direct imaging of a natural-satellite transit from the surface of another planet. The sun is ≈ 20 pixels in diameter on the images, with Mars I and II having apparent diameters of ≈ 10-13 and 2 pixels, respectively. Mars II was already in transit when the first image was obtained on Mar. 4.12773 UT, was approximately at mid-transit on the Mar. 4.12785 image, and was between third and fourth contacts on the Mar. 4.12819 image. Mars I was possibly near first contact on the image of Mar. 7.11556, was near mid-transit on Mar. 7.11567, and was near fourth contact on Mar. 7.11579.

2004 March 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green